# Rethinking of Feature Interaction for Multi-task Learning on Dense Prediction

### v1

@@ 3.1 Overview
The overview of the proposed BFCI is shown in Fig.2, which mainly consists of TPP for task-pattern disentanglement, BFE for bridge feature extraction, and TFR for task-specific feature refinement. In the early decoding stage, we gain multi-scale task-generic features $\mathbf{S}=\left\{S_{i}\mid i=0,1,2,3\right\}$ from the shared image backbone, which can be ConvNets (e.g. HRNet, ResNet) or Vision Transformers (e.g. ViT). The task-generic features are firstly partitioned into patch tokens and embedded into the decoding dimensions. Simultaneously, a set of preliminary decoders with deep supervision are implemented to produce task-specific features $\mathbf{P}=\left\{P^{j}_{i}\mid i=0,1,2,3;\quad j=0,1,\cdots,T\right\}$, where $T$ represents the total number of tasks. The TPP is applied before the feature is fed into initial prediction heads to avoid task-pattern entanglement. Subsequently, both task-generic and -specific features are fed into BFE to produce the multi-scale bridge features $\mathbf{S^{\prime}}=\left\{S^{\prime}_{i}\mid i=0,1,2,3\right\}$. After the bridge features are formed, they are in turn used to fertilize task-specific features by TFR which transfers helpful representations from bridge features to task-specific features. Both BFE and TFR are applied on multiple scales, the task-specific features are up-sampled by shared up-sampling layers at each scale. Finally, the refined multi-scale task-specific features are aggregated and fed into prediction heads to make final predictions. The detailed structures will be described in the next sections.
Figure 4: Left: Visualization of attention patterns of two tasks in TPP. The depth feature focuses more on the surface and edge features with geometry information, while semantic feature focuses on objects with various semantic information, like cabinets and floors. Right: Visualization of the task-specific features before and after TPP. The last scale of encoder outputs shows the task-pattern-entanglement issue, which behaves significantly differently from task labels and the representations they contain are implicit and ambiguous. With TPP, the task-specific features can produce well-decoupled representations more similar to their ground-truth distributions (like the highlighted floor area in semantic segmentation and door area in depth estimation).
@@ 3.2 Bridge Feature Extraction
$\begin{array}[]{c}\widetilde{S_{i}}=\operatorname{Flat}(\operatorname{Conv}(S_{i}))\\ \widetilde{P_{i}^{j}}=\operatorname{Flat}(\operatorname{Pool}(P_{i}^{j})),\quad j=0,1,\cdots,T\end{array}$ (1)
@@
$\begin{array}[]{c}Q=\operatorname{LP}(\widetilde{S_{i}})\\ K=\operatorname{LP}(\operatorname{Concat}(\widetilde{P_{i}^{0}},\widetilde{P_{i}^{1}},\cdots,\widetilde{P_{i}^{T}})\\ V=\operatorname{LP}(\operatorname{Concat}(\widetilde{P_{i}^{0}},\widetilde{P_{i}^{1}},\cdots,\widetilde{P_{i}^{T}})\end{array}$ (2)
where $Q\in\mathbb{R}^{C_{a}\times\left(\frac{H_{s}W_{s}}{{d}^{2}}\right)}$ and $K,V\in\mathbb{R}^{C_{a}\times\left(\frac{TH_{s}W_{s}}{{l}^{2}}\right)}$, $C_{a}$ is the channel dimension for attention. Then, we conduct a standard cross-attention:
$A=\frac{Q\times K^{T}}{\sqrt{C_{a}}},\quad A\in\mathbb{R}^{\frac{H_{s}W_{s}}{d^{2}}\times\frac{TH_{p}W_{p}}{l^{2}}}$ (3)
@@
$x=\operatorname{Softmax}(A)\times V,\quad x\in\mathbb{R}^{C_{s}\times\left(\frac{H_{s}W_{s}}{{d}^{2}}\right)}$ (4)
then $x$ is reshaped and up-sampled into ${C_{a}\times\left({H_{s}W_{s}}\right)}$ and fed into a feed-forward network which is composed of a layer normalization and an MLP. The final output is reshaped into ${C_{a}\times H_{s}\times W_{s}}$, which is the bridge feature we denote as $S^{\prime}_{i}$ at scale $i$. By querying all of the task-specific features globally, the task-generic feature gains discriminative high-level representations by selecting the task-specific pixels that have the highest response to the task-generic features in $A$, and the low-level representations are simultaneously reserved by residual paths. Thus, the extracted bridge feature meets the need as a medium for multi-task dense prediction. Different from directly conducting pixel-wise addition of task-generic and -specific features, which might cause unexpected conflicts if there is task inconsistency exists in a certain position, the global querying ensures only the highest responding task-specific pixel is selected and avoids the potential task conflicts.
3.3 Task Pattern Propagation
For multi-scale task-generic features, high-level features with smaller scales contain richer semantic and more contextual information and usually guide the formation of predictions in a top-down way. Meanwhile, different dense prediction tasks usually share a lot of similarities in low-level representations, but significantly differ in high-level representations. Thus, for the task-generic features produced by the final layers of the shared backbone, the high-level representations from different tasks are usually entangled severely, making it difficult to extract discriminative representations for each task and subsequently conduct interactions. We name this phenomenon the task-pattern-entanglement issue. As shown in Figure 4 (left), we visualize the regions with soft attention scores on semantic segmentation and depth estimation respectively, which show great differences that depth attention focuses more on spatial and geometry information while semantic attention focuses more on different objects and contexts. In Figure 4 (right), the patterns in the encoder shared output are implicit and entangled, which do not clearly reflect the characteristics of either task.
To tackle the task-pattern-entanglement issue, we propose to conduct Task Pattern Propagation (TPP) that jointly learns and assigns different task patterns to corresponding tasks to assist in the formation of task-specific features. Although learning and propagating task patterns have been discussed in previous multi-task learning works [53, 55], the entanglement issues at the pattern level remain unexplored and unsolved, and our proposed TPP first time targeting this issue from the task pattern level. As shown in Fig. 5, firstly, for each task-specific feature maps ${P_{3}^{j}}$ at scale $3$, we generate $T$ groups of query $Q_{j}$, key $K_{j}$, and value $V_{j}$, where $j=1,2,\cdots,T$. The generation process is similar to Sec. 3.2. Then $T$ groups of task-specific attentions $A_{j}$ are calculated by dot-product, which contains patterns of each task as follows:
$A_{j}=\frac{Q_{j}\times K_{j}^{T}}{\sqrt{C_{a}}}$ (5)
where $C_{a}$ denotes the dimension for attention space. Subsequently, attention maps from all tasks $A_{j},j=1,2,\cdots,T$ are concatenated and then squeezed by $1\times 1$ convolution in a common space to align the dimension and share task patterns. Later the task patterns in attention maps are propagated by dot-product with all of the task values:
$\begin{array}[]{c}\bar{A}=\operatorname{Conv_{1\times 1}}\left(\operatorname{Concat}\left(A_{1},A_{2},\cdots,A_{T}\right)\right)\\ x_{j}=\operatorname{Softmax}(\bar{A})\times V_{j},\quad j=1,2,\cdots,T\end{array}$ (6)
Finally, $x_{j}$ are processed by a FFN which has a similar structure in BFE to enhance the learned task-specific high-level representations. Since different tasks have various pattern distributions, it is necessary to perform a pattern propagation to share the attention space for multiple tasks. Empirically, the irrelevant pixels will be repressed during the matching to avoid the negative feature from transferring, which avoids the unexpected entanglement caused by task inconsistencies. With TPP, we extract explicit task patterns by self-attention and conduct propagation to avoid potential entanglements and achieve a clear decoupling effect of task-specific features, which also benefits the subsequent feature interactions of BFE and TFR. As shown in Fig. 4 (Right), TPP can significantly strengthen the task feature by producing more discriminative high-level representations.
@@ 3.4 Task Feature Refiner
$\begin{array}[]{c}x^{(k)}=\operatorname{TFRlayer}^{(k)}\left(\operatorname{Concat}\left(S_{i}^{\prime},x^{(k-1)}\right)\right),\\ j=0,1,\cdots,T;\quad k=1,2,\cdots,M;\quad x^{(0)}=P_{i}^{j}\end{array}$ (7)
where $M$ represents the total number of layers in TFR, stacking these TFR layers forms the entire TFR module. During this process, the task-specific features $P_{i}^{j}$ undergoes progressive refinement through these layers, continuously guided by the bridge features, obtaining abundant task representations. This ultimately generates high-quality task-specific features, laying a solid foundation for the generation of final task predictions.
Additionally, our TFR layers can be flexibly configured. In the simplest case, we can use just one convolutional layer to align the concatenated feature channels and generate predictions. However, this is certainly insufficient, as a single convolutional layer lacks the necessary capacity for nonlinear transformations. Alternatively, Transformer layers can be employed for global relationship modeling, such as InvPT [50]. Yet, experimental findings reveal that the complex quadratic relationship modeling of the Transformer does not significantly improve the quality of task features. This is partly because the scales of the feature maps are not large at this stage, and global modeling doesn’t offer significant advantages. Furthermore, spatial local similarity serves as a more crucial prior for dense predictions. Hence, in our method, we utilize depth-wise separable dilated convolutions to form our fundamental layer. We stack dilated convolutions of varying sizes to ensure sufficient receptive fields and avoid the grid-effect associated with dilation. Simultaneously, the local connections brought by convolutions are suitable for extracting vital local information from the bridge features to generate the final predictions.
Figure 6: As visualized above, we compare with SOTA works (ATRC[3] and InvPT [50]) on NYUD-v2 4 tasks. Ours BFCI produces clearly more precise predictions on each task.
@@ 4.1 Experimental Setup
Datasets: We conduct experiments on two benchmark datasets: NYUD-v2 [34] and PASCAL Context [7]. The NYUD-v2 dataset is mainly used for indoor scene segmentation and depth estimation in MTL works. The dataset contains 1449 images. Following the standard settings, we use 795 images for training which are randomly selected, and the rest are for testing. The PASCAL dataset is a popular benchmark for many dense prediction tasks and we use the split from PASCAL Context, which contains 10103 images, where 4998 images are randomly selected for training and the rest are for testing. We choose several subsets of tasks on both datasets to make a more explicit comparison, including semantic segmentation (Seg.), human parts segmentation (H.Parts), depth eatimation (Dep.), surface normal estimation (Norm.), saliency estimation (Sal.) and edge detection (Edge.).
@@
BFCI (ours) 56.57 0.4655 17.29 80.02 BFCI (ours) 79.89 71.33 85.64 13.38 73.24
TABLE II: Comparison with SOTA works on different backbones.
@@
BFCI 55.51 0.4930 18.47 78.22 BFCI 80.64 70.06 84.64 13.82 72.96
TABLE III: Analysis of model sizes and computation cost.
Models FLOPs Params Seg. H.Parts Sal. Norm. Edge.
(G) (M) mIoU mIoU maxF mErr odsF
PAD-Net [48] w. ViT-large 773 330 78.01 67.12 79.21 14.37 72.60
MTI-Net [38] w. ViT-large 774 851 78.31 67.40 84.75 14.67 73.00
ATRC [3] w. ViT-large 871 340 77.11 66.84 81.20 14.23 72.10
InvPT [50] w. ViT-large 669 423 79.03 67.61 84.81 14.15 73.00
BFCI w. ViT-base 412 230 77.98 68.19 85.06 13.48 72.70
BFCI w. ViT-large 645 477 80.64 70.06 84.64 13.82 72.96
BFCI w. InternImage-B 406 248 78.07 69.03 85.29 13.44 73.10
BFCI w. InternImage-L 527 420 79.89 71.33 85.64 13.38 73.24
@@
$\Delta_{{\tau}_{j}}=(-1)^{l_{j}}\left(M_{m,j}-M_{s,j}\right)/M_{s,j}$ (8)
@@
$\Delta_{MTL}=\frac{1}{N}\sum_{j=1}^{N}\Delta_{{\tau}_{j}}$ (9)
Implementation Details: We conduct our experiments on Pytorch [31] with one NVIDIA Tesla V100 GPU. The models used for different evaluation experiments were trained for $40,000$ iterations on both datasets with a batch size of 6. We employed various backbone networks to comprehensively validate our approach. These include classic single-scale CNN encoders like the ResNet [16] series with dilated convolutions, single-scale Transformer encoders like the ViT [10] series, multi-scale dense prediction network HRNet [35] series, and multi-scale deformable convolution series InternImage [41]. For models using ViT and ResNet-50 backbones, we utilized the Adam optimizer with a learning rate of $2\times 10^{-5}$ and a weight decay rate of $1\times 10^{-6}$. For models using HRNet backbones, we utilized the SGD optimizer with a learning rate of $0.01$ and a weight decay rate of $5\times 10^{-4}$, and the momentum weight is $0.9$. And for models using InternImage backbones, we utilized the AdamW optimizer with a learning rate of $6\times 10^{-5}$and $2\times 10^{-5}$ for the base and the large model respectively, and a weight decay rate of $0.05$, the $\beta$s for the optimizers are $(0.9,0.999)$. A polynomial learning rate decay scheduler was used. We maintained alignment with the settings from [50], first converting the output resolution and channel count of the backbone network into a multi-scale feature pyramid. The last three layers were taken as the multi-scale input for BFCI. Correspondingly, BFE and TFR were applied only to features from these last three scales, ensuring a favorable trade-off between performance and inference speed.
@@
Figure 7: As visualized above, we compare with SOTA works (ATRC[3] and InvPT [50]) on PASCAL Context 5 tasks. Ours BFCI produces clearly more precise predictions on each task.

### v2

@@ 3.1 Overview
The overview of the proposed BridgeNet is shown in Fig.2, which mainly consists of TPP for task-pattern disentanglement, BFE for bridge feature extraction, and TFR for task-specific feature refinement. In the early decoding stage, we gain multi-scale task-generic features $\mathbf{S}=\left\{S_{i}\mid i=0,1,2,3\right\}$ from the shared image backbone, which can be ConvNets (e.g. HRNet, ResNet) or Vision Transformers (e.g. ViT). The task-generic features are firstly partitioned into patch tokens and embedded into the decoding dimensions. Simultaneously, a set of preliminary decoders with deep supervision are implemented to produce task-specific features $\mathbf{P}=\left\{P^{j}_{i}\mid i=0,1,2,3;\quad j=0,1,\cdots,T\right\}$, where $T$ represents the total number of tasks. The TPP is applied before the feature is fed into initial prediction heads to avoid task-pattern entanglement. Subsequently, both task-generic and -specific features are fed into BFE to produce the multi-scale bridge features $\mathbf{S^{\prime}}=\left\{S^{\prime}_{i}\mid i=0,1,2,3\right\}$. After the bridge features are formed, they are in turn used to fertilize task-specific features by TFR which transfers helpful representations from bridge features to task-specific features. Both BFE and TFR are applied on multiple scales, the task-specific features are up-sampled by shared up-sampling layers at each scale. Finally, the refined multi-scale task-specific features are aggregated and fed into prediction heads to make final predictions. The detailed structures will be described in the next sections.
3.2 Task Pattern Propagation
For multi-scale task-generic features, features with smaller scales contain richer high-level semantic and more contextual information and usually guide the formation of predictions in a top-down way. Meanwhile, different dense prediction tasks usually share a lot of similarities in low-level representations, but significantly differ in high-level representations. Thus, for the task-generic features produced by the final layers of the shared backbone, the high-level representations from different tasks are usually entangled severely, making it difficult to extract discriminative representations for each task and subsequently conduct interactions. We name this phenomenon the task-pattern-entanglement issue. As shown in Figure 3 (left), we visualize the regions with soft attention scores on semantic segmentation and depth estimation respectively, which show great differences that semantic attention focuses more on different objects and contexts while depth attention focuses more on spatial and geometry information. In Figure 3 (right), the patterns in the encoder shared outputs are implicit and entangled, which do not clearly reflect the distributions of either task. Thus, task-specific features directly decomposed from them are struggling with the lack of discriminative semantics, which leads to low-quality interaction participants for the following interactions and eventually brings negative effects to interaction outputs.
To tackle the task-pattern-entanglement issue, we propose to conduct Task Pattern Propagation (TPP) that jointly learns and propagates different task patterns to corresponding tasks to assist in the formation of task-specific features. Although learning and propagating task patterns have been discussed in previous multi-task learning works [55, 57], the entanglement issues at the pattern level remain unexplored and unsolved, and our proposed TPP first time targeting this issue from the task pattern level. As shown in Fig. 4, firstly, for every task-specific feature maps ${P_{3}^{j}}$ at scale $3$, where $j=1,2,\cdots,T$, we correspondingly generate $T$ sets of query $Q_{j}$, key $K_{j}$, and value $V_{j}$ by linear projection for each task. Then task-specific attentions $A_{j},j=1,2,\cdots,T$ are calculated by dot-product accordingly, which contains attention patterns of each task as follows:
$A_{j}=\frac{Q_{j}\times K_{j}^{\top}}{\sqrt{C_{a}}},$ (1)
where $C_{a}$ denotes the dimension for attention space. Subsequently, attention maps from all tasks $A_{j}$ are concatenated and then squeezed by $1\times 1$ convolution in a common space to align the dimension and share task patterns. Later the task patterns in attention maps are propagated by dot-product with all of the task values:
$\begin{array}[]{c}\bar{A}=\operatorname{Conv_{1\times 1}}\left(\operatorname{Concat}\left(A_{1},A_{2},\cdots,A_{T}\right)\right),\\ x_{j}={\operatorname{Softmax_{dim=-1}}}(\bar{A})\times V_{j},\quad j=1,2,\cdots,T.\end{array}$ (2)
where $\operatorname{Softmax_{dim=-1}(\cdot)}$ represents the softmax function applied to the last feature dimension.
Finally, $x_{j}$ are processed by an FFN which has a similar structure in BFE to enhance the learned task-specific high-level representations. Since different tasks have various pattern distributions, it is necessary to perform a pattern propagation to share the attention space for multiple tasks. Empirically, the irrelevant pixels will be repressed during the matching to avoid the negative feature from transferring, which avoids the unexpected entanglement caused by task inconsistencies. With TPP, we extract explicit task patterns by self-attention and conduct propagation to avoid potential entanglements and achieve a clear decoupling effect of task-specific features, which also benefits the subsequent feature interactions of BFE and TFR. As shown in Fig. 3 (right), TPP can significantly strengthen the task patterns by producing more discriminative high-level semantics, like the highlighted floor area in semantic segmentation and door area in depth estimation, which boosts the subsequent interactions by providing high-quality participants.
@@ 3.3 Bridge Feature Extraction
$\begin{array}[]{c}\widetilde{S_{i}}=\operatorname{Flat}(\operatorname{Conv}(S_{i})),\\ \widetilde{P_{i}^{j}}=\operatorname{Flat}(\operatorname{Pool}(P_{i}^{j})),\quad j=0,1,\cdots,T,\end{array}$ (3)
@@
$\begin{array}[]{c}Q=\operatorname{LP}(\widetilde{S_{i}})^{\top},\\ K=\operatorname{LP}(\operatorname{Concat}(\widetilde{P_{i}^{0}},\widetilde{P_{i}^{1}},\cdots,\widetilde{P_{i}^{T}})^{\top},\\ V=\operatorname{LP}(\operatorname{Concat}(\widetilde{P_{i}^{0}},\widetilde{P_{i}^{1}},\cdots,\widetilde{P_{i}^{T}})^{\top},\end{array}$ (4)
where $Q\in\mathbb{R}^{\left(\frac{H_{s}W_{s}}{{d}^{2}}\right)\times C_{a}}$ and $K,V\in\mathbb{R}^{\left(\frac{TH_{p}W_{p}}{{l}^{2}}\right)\times C_{a}}$, $C_{a}$ is the channel dimension for attention. Then, we conduct a standard cross-attention:
$A=\frac{Q\times K^{\top}}{\sqrt{C_{a}}},\quad A\in\mathbb{R}^{\frac{H_{s}W_{s}}{d^{2}}\times\frac{TH_{p}W_{p}}{l^{2}}},$ (5)
@@
$x={\operatorname{Softmax_{dim=-1}}}(A)\times V,\quad x\in\mathbb{R}^{\left(\frac{H_{s}W_{s}}{{d}^{2}}\right)\times C_{a}},$ (6)
Then $x$ is transposed, reshaped, and up-sampled into ${C_{a}\times\left({H_{s}W_{s}}\right)}$ and fed into a feed-forward network which is composed of a layer normalization and an MLP. The final output is reshaped into ${C_{a}\times H_{s}\times W_{s}}$, which is the bridge feature we denote as $S^{\prime}_{i}$ at scale $i$. By querying all of the task-specific features globally, the task-generic feature gains discriminative high-level representations by selecting the task-specific pixels that have the highest response to the task-generic features in $A$, and the low-level representations are simultaneously reserved by residual paths. Thus, the extracted bridge feature meets the need as a medium for multi-task dense prediction. Different from directly conducting pixel-wise addition of task-generic and -specific features, which might cause unexpected conflicts if there is task inconsistency exists in a certain position, the global querying ensures only the highest responding task-specific pixel is selected and avoids the potential task conflicts.
@@ 3.4 Task Feature Refiner
$\begin{array}[]{c}x^{(k)}=\operatorname{TFRlayer}^{(k)}\left(\operatorname{Concat}\left(S_{i}^{\prime},x^{(k-1)}\right)\right),\\ j=0,1,\cdots,T;\quad k=1,2,\cdots,M;\quad x^{(0)}=P_{i}^{j},\end{array}$ (7)
where $M$ represents the total number of layers in TFR, stacking these TFR layers forms the entire TFR module. During this process, the task-specific features $P_{i}^{j}$ undergo progressive refinement through these layers, continuously guided by the bridge features, obtaining abundant task representations. This ultimately generates high-quality task-specific features, laying a solid foundation for the generation of final task predictions.
Additionally, our TFR layers can be flexibly configured. In the simplest case, we can use just one convolutional layer to align the concatenated feature channels and generate predictions. However, this is certainly insufficient, as a single convolutional layer lacks the necessary capacity for nonlinear transformations. Alternatively, Transformer layers can be employed for global relationship modeling, such as InvPT [51]. Yet, experimental findings reveal that the complex quadratic relationship modeling of the Transformer does not significantly improve the quality of task features. This is partly because the scales of the feature maps are not large at this stage, and global modeling doesn’t offer significant advantages. Furthermore, spatial local similarity serves as a more crucial prior for dense predictions. Hence, in our method, we utilize depth-wise separable dilated convolutions to form our fundamental layer. We stack dilated convolutions of varying sizes to ensure sufficient receptive fields and avoid the grid effect associated with dilation. Simultaneously, the local connections brought by convolutions are suitable for extracting vital local information from the bridge features to generate the final predictions.
Figure 6: As visualized above, we compare with SOTA works (ATRC[3] and InvPT [51]) on NYUD-v2 4 tasks. Our BridgeNet clearly produces more precise predictions on each task. To avoid cherry-picking, the input image is selected from the sample examples provided in [51] paper.
@@ 4.1 Experimental Setup
Datasets: We conduct experiments on three benchmark datasets: NYUD-v2 [35], Cityscapes [9] and PASCAL Context [7]. NYUD-v2 dataset is mainly used for indoor scene segmentation and depth estimation in MTL works. The dataset contains 1449 images. Following the standard settings in [38], we use 795 images for training which are randomly selected, and the rest are for testing. Cityscapes [9] contains 2975 and 500 high-resolution street-view images captured from different European countries for training and testing respectively. These images are annotated with fine 19-class semantic labels for segmentation and disparity map for depth estimation. To speed up the convergence and train all models in comparisons with controllable computational cost overhead, we down-sample the images with their annotations to $384\times 768$ resolution. PASCAL is a popular benchmark for many dense prediction tasks and we use the split from PASCAL Context, which contains 10103 images, where 4998 images are randomly selected for training and the rest are for testing [38]. We choose several subsets of tasks on both datasets to make a more explicit comparison, including semantic segmentation (Seg.), human parts segmentation (H.Parts), depth eatimation (Dep.), surface normal estimation (Norm.), saliency estimation (Sal.) and edge detection (Edge.).
@@
BridgeNet (ours) 56.57 0.4655 17.29 80.02 BridgeNet (ours) 79.89 71.33 85.64 13.38 73.24
TABLE II: Comparison with SOTA works on Cityscapes.
       Cityscapes
       Models        Seg.        Dep.
       mIoU        rmse
       Cross-Stitch [30]        85.31        3.422
       NDDR-CNN [15]        84.73        3.415
       MTAN [24]        90.20        3.442
       PAD-Net [49]        90.13        3.496
       MTI-Net [39]        91.50        2.680
       InvPT [51]        91.19        2.631
       BridgeNet (ours)        92.61        2.606
Figure 7: As visualized above, we compare with SOTA works (including PAD-Net [49], MTAN [24], MTI-Net [39] and InvPT [51]) on Cityscapes.
TABLE III: Comparison with SOTA works on different backbones.
@@
BridgeNet 55.51 0.4930 18.47 78.22 BridgeNet 80.64 70.06 84.64 13.82 72.96
TABLE IV: Analysis of model sizes and cost. Bold and underlined fonts represent the first and second best results respectively.
Models FLOPs Params Latency Memory Seg. H.Parts Sal. Norm. Edge.
(GB) (M) (s) (GB) mIoU mIoU maxF mErr odsF
PAD-Net [49] w. ViT-large 773 330 0.043 3.096 78.01 67.12 79.21 14.37 72.60
MTI-Net [39] w. ViT-large 774 851 0.087 7.642 78.31 67.40 84.75 14.67 73.00
ATRC [3] w. ViT-large 871 340 0.098 7.933 77.11 66.84 81.20 14.23 72.10
InvPT [51] w. ViT-large 669 423 0.060 3.510 79.03 67.61 84.81 14.15 73.00
BridgeNet w. ViT-base 412 230 0.090 2.710 77.98 68.19 85.06 13.48 72.70
BridgeNet w. ViT-large 645 477 0.144 3.718 80.64 70.06 84.64 13.82 72.96
BridgeNet w. InternImage-B 406 248 0.130 2.774 78.07 69.03 85.29 13.44 73.10
BridgeNet w. InternImage-L 527 420 0.171 3.494 79.89 71.33 85.64 13.38 73.24
@@
$\Delta_{{\tau}_{j}}=(-1)^{l_{j}}\left(M_{m,j}-M_{s,j}\right)/M_{s,j},$ (8)
@@
$\Delta_{MTL}=\frac{1}{N}\sum_{j=1}^{N}\Delta_{{\tau}_{j}},$ (9)
Implementation Details: We conduct our experiments on Pytorch [32] with one NVIDIA Tesla V100 GPU. The models used for different evaluation experiments were trained for $40,000$ iterations on both datasets with a batch size of 6. We employed various backbone networks to comprehensively validate our approach. These include classic single-scale CNN encoders like the ResNet [17] series with dilated convolutions, single-scale Transformer encoders like the ViT [11] series, multi-scale dense prediction network HRNet [36] series, and multi-scale deformable convolution series InternImage [42]. For models using ViT and ResNet-50 backbones, we utilized the Adam optimizer with a learning rate of $2\times 10^{-5}$ and a weight decay rate of $1\times 10^{-6}$. For models using HRNet backbones, we utilized the SGD optimizer with a learning rate of $0.01$ and a weight decay rate of $5\times 10^{-4}$, and the momentum weight is $0.9$. And for models using InternImage backbones, we utilized the AdamW optimizer with a learning rate of $6\times 10^{-5}$and $2\times 10^{-5}$ for the base and the large model respectively, and a weight decay rate of $0.05$, the $\beta$ for the optimizers are $(0.9,0.999)$. A polynomial learning rate decay scheduler was used.
We maintained alignment with the settings from [51]. For single-scale image backbones, we first extract the intermediate backbone features at specified depths, then convert their spatial and channel dimension by convolution layers to build a multi-scale feature pyramid. The last three layers were taken as the multi-scale input for BridgeNet. Correspondingly, BFE and TFR were applied only to features from these last three scales, ensuring a favorable trade-off between performance and inference speed. For multi-scale image backbones, they directly produce a feature pyramid which is acceptable for BridgeNet.
@@
Figure 8: As visualized above, we compare with SOTA works (ATRC[3] and InvPT [51]) on PASCAL Context 5 tasks. Our BridgeNet clearly produces more precise predictions on each task. We also use the sample example image provided in [51] paper for comparisons.